# Efficient Statistical Estimation for Sequential Adaptive Experiments with Implications for Adaptive Designs

### v1

@@ 8 Application and Generalization
8.1 Application in Independent Settings
Our framework is directly applicable for independent setups. For example, consider estimating causal estimands from data collected across multiple sites where the site index does not affect the primary outcome but introduces heterogeneity in treatment distribution across sites. Suppose the full data of interest is $X=(W,Y(a))\sim P^{F}\in\mathcal{M}^{F}$, where $W$ denotes the baseline; $Y(a)$ is the counterfactual outcome with treatment level $a\in\mathcal{A}$. Let the average treatment effect $\Psi^{F}(P^{F})=E_{P^{F}}[Y(1)-Y(0)]$ as the causal estimand of interest. Suppose one observe $n$ independent observations $O_{1},\cdots,O_{n}$ with $O_{i}=(S_{i},W_{i},A_{i},Y_{i})$, where $W_{i}\sim Q_{W}$ and $Y_{i}\sim Q_{Y}(\cdot|A_{i},W_{i})$ for $Q_{W}\in\mathcal{Q}_{W}$ and $Q_{Y}\in\mathcal{Q}_{Y}$. For every unit $i$, $S_{i}\in\{1,\dots,J\}$ is the site index that determines the treatment distribution of $A_{i}\sim g_{i}$ with $g_{i}=g_{s_{i}}$, as each site has its own treatment mechanisms $g_{j}:\mathcal{W}\times\mathcal{A}\to(0,1)$ for $j\in[J]$. $\Psi^{F}(P^{F})$ can be identified by the target parameter $\psi(Q)=E[\bar{Q}(1,W)-\bar{Q}(0,W)]$ under ignorability and positivity assumptions, where $\bar{Q}(A,W):=E[Y|A,W]$. Note that each site does not impact the outcome given treatment and baseline covariates.
Define $\bar{g}_{n}:=\frac{1}{n}\sum_{i=1}^{n}g_{i}$. With the convergence of $\bar{g}_{n}$ to a fixed $\bar{g}_{\infty}$ (e.g. $\bar{g}_{\infty}=\sum_{j\in J}p_{s}(j)g_{j}$ for a probability mass function $p_{s}$ of the site index), one can construct a ADL-TMLE by solving the score equation $\frac{1}{n}\sum_{i=1}^{n}\frac{2A_{i}-1}{\bar{g}_{n}(A_{i}|W_{i})}[Y_{i}-\bar{Q}_{n}^{*}(A_{i},W_{i})]=0$. This yields an asymptotic normal and efficient estimator as established in the general framework above for adaptive design setups. This also extends to other settings where the site index $S$ is replaced by other variables that do not directly affect the outcome given treatment and other baseline covariates, but only influence the treatment assignment probabilities.
8.2 A General Framework
We present a general estimation and design framework for various causal estimands in different setups including those with longitudinal data structure.
Data, Model and Target Parameter. Consider a longitudinal data structure $O=\left(L(0),A(0),L(1),A(1),\cdots,L(K),A(K),L(K+1)\equiv Y\right)$. Let
@@ 8.2 A General Framework
be a model with a known $g$, where $q(o)=\prod_{t=0}^{K+1}q_{t}(o)$ and $g(o)=\prod_{t=0}^{k}g_{t}(o)$, which corresponds to an i.i.d. likelihood with each $L(t)\sim q_{t}$ and $A(t)\sim g_{t}$. Let $g^{*}\in\mathcal{G}^{*}$ denote a fixed function that defines the causal parameter of interest as $\Psi^{F}(Q)=EY_{g^{*}}=\int_{o}y\,q(o)\,g^{*}(o)\,d\mu(o)$. This parameter only depends on the $Q$-component of the likelihood.
@@
$p^{n}_{q,\mathbf{g}_{1:n}}(\bar{o}(n))=\prod_{i=1}^{n}q(o_{i})\,g_{i}(o_{i})=\prod_{i=1}^{n}\prod_{t=1}^{K+1}q_{t}(o_{i})\prod_{t=1}^{K}g_{i,t}(o_{i}).$
@@
Efficient Estimator. Define $\bar{g}_{n}=\frac{1}{n}\sum_{i=1}^{n}g_{i}$ as the average of the whole product $\prod_{t}g_{t,i}$ across $i=1,\cdots,n$. Let $D(Q,g)(O_{i})$ to denote the efficient influence curve in the i.i.d. model $\mathcal{M}(g)$. To estimate $EY_{g^{*}}$ from data generated by the model $\mathcal{M}^{n}$, it is useful to consider a reference i.i.d. model $\mathcal{M}(\bar{g}_{n})$ with the average design $\bar{g}_{n}$. In that i.i.d. model, the efficient influence curve for the parameter $\Psi(Q)$ is given by $D(Q,\bar{g}_{n})(O)$. Importantly, any TMLE constructed by solving $n^{-1}\sum_{i=1}^{n}D(Q,\bar{g}_{n})(O_{i})$ under the i.i.d. model with $\bar{g}_{n}$ is also an efficient estimator for the original model $\mathcal{M}^{n}$ for the observed data, with its asymptotic normality supported by Martingale theories and its efficiency provided by the convolution theorem, as shown in Section 4 and 5. This unified estimation framework applies not only to adaptive design settings discussed above, but also to other contexts with longitudinal data structure with dynamic interventions (Robins 1986, van der Laan & Gruber 2012) and time-series structures (van der Laan & Malenica 2018, Malenica et al. 2021). For the design problem, the self-calibrating $\bar{g}$-driven adaptive design implied by this efficient estimation approach directly mirrors the construction in Section 6.
@@ 9 Conclusions
In this work, we propose an efficient framework for estimating a wide class of causal estimands from sequential adaptive experiments based on ADL-TMLE. We establish its asymptotic normality and semiparametric efficiency, present relaxed positivity and stabilization assumptions for adaptive experiments, and show its superior finite-sample performance in variance reduction. Motivated by efficiency results, we further propose a $\bar{g}$-driven adaptive design that steers the average design toward the estimated optimal design in a stable self-calibrating manner, converges to the oracle design more quickly, and improves estimation efficiency compared with the standard approach. Finally, we provide a general estimation and design framework applicable to diverse settings including those with longitudinal structures. Future research could explore extensions to other related settings and broader areas of applications.
@@
Combining with the fact that ADL-TMLE solves $\frac{1}{n}\sum_{i=1}^{n}D(Q_{n}^{*},\bar{g}_{n})(O_{i})=0$, we have $\psi(Q_{n}^{*})-\psi(Q_{0})=\frac{1}{n}\sum_{i=1}^{n}[D(Q_{n}^{*},\bar{g}_{n})(O_{i})-P_{Q_{0},\bar{g}_{n}}D(Q_{n}^{*},\bar{g}_{n})$]. By adding and subtracting $D(Q_{\infty},\bar{g}_{n})(O_{i})$, $D(Q_{\infty},\bar{g}_{\infty})(O_{i})$, and their conditional expectations given $\bar{O}(i-1)$, we obtain the desired decomposition. ∎
@@
$=$ $\Delta_{n,h}-\frac{1}{2}\|h\|_{\mathcal{H}(Q,\bar{g}_{\infty})}^{2}$
where we denote $\Delta_{n,h}:=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}h(O_{i})+o_{P^{n}_{Q}}(1)$, with the first term a zero-mean martingale. The second equation is based on the convergence of conditional variance of the zero-mean martingale according to Theorem 2.23 of Hall & Heyde (2014) as $E[h^{2}(O_{i})|\bar{O}(i-1)]$ is bounded. The third equation is based on the dependence of $g_{i}$ on the past $\bar{O}(i-1)$ in our adaptive design setup and the definition of $\|h\|_{\mathcal{H}(Q,\bar{g}_{n})}^{2}$. The last equation is based on the convergence of the adaptive design in Assumption 10 and results in $\Delta_{n,h}\overset{d}{\to}\Delta_{h}\sim N(0,\|h\|_{\mathcal{H}(Q,\bar{g}_{\infty})}^{2})$ by Martingale central limit theorem. ∎
@@
With Assumption 10, one can obtain $log\frac{dP^{n}_{Q_{\epsilon_{n}}^{h}}(O_{1},\cdots,O_{n})}{dP^{n}_{Q}(O_{1},\cdots,O_{n})}=\Delta_{n,h}-\frac{1}{2}\|h\|_{\mathcal{H}(Q_{0},\bar{g}_{\infty})}^{2}$ with $\Delta_{n,h}\overset{d}{\to}\Delta_{h}\sim N(0,\|h\|_{\mathcal{H}(Q_{0},\bar{g}_{\infty})}^{2})$ by Theorem 7. Then, by Le Cam’s third lemma (e.g. see van der Vaart et al. (1996)), we have that
@@
By convolution theorem (Theorem 8), the asymptotic variance of $\sqrt{n}(T_{n}-\psi(Q_{0}))$ for regular estimators $T_{n}$ is no less than $\|\dot{\psi}^{*}(Q_{0})\|_{\mathcal{H}(Q_{0},\bar{g}_{\infty})}^{2}$, which equals $P_{Q_{0},\bar{g}_{\infty}}D(Q_{0},\bar{g}_{\infty})^{2}$ and is achieved by the asymptotic variance of $\psi(Q_{n}^{*})$. Therefore, $\psi(Q_{n}^{*})$ is efficient. ∎
@@
The data generating distribution is as follows: $\bar{Q}_{0}(1,W)=25+10/\left[1+e^{-(2W+1)}\right]$, $\bar{Q}_{0}(0,W)=20+17.5/\left[1+e^{-(W+0.1)}\right]$, $Var_{Q_{0}}[Y|A=1,W]=1+3W^{3}$, $Var_{Q_{0}}[Y|A=0,W]=1+1.5\times(3-W)^{3}.$
For benefit-driven design, we guide treatment randomization probabilities toward optimal personalized treatment using baseline covariates through estimating the conditional average treatment effect (CATE). Define the true CATE function as $B_{0}:\mathcal{W}\to[-1,1]$, with $B_{0}(W)=\bar{Q}_{0}(1,W)-\bar{Q}_{0}(0,W)$. For a newly enrolled unit $i$ with baseline covariates $w_{i}$, we use $B_{n}(w_{i})$ to denote the estimate of $B_{0}(w_{i})$ based on the current data $\bar{O}(i-1)$. We estimate CATE functions by first transforming outcomes into doubly robust pseudo-outcomes (Rubin & van der Laan 2005, van der Laan & Rubin 2006, van der Laan & Luedtke 2015, Kennedy 2023), then regressing them on baseline covariates using Super Learner (van der Laan et al. 2007). This estimated CATE guides the treatment randomization probabilities toward the estimated optimal personalized treatment $I(B_{n}(w_{i})>0)$ through a function $\Gamma_{\nu,b}$ introduced by Chambaz et al. (2017):
@@
Here, $B_{n}(w_{i})$ is used as input to $\Gamma_{\nu,b}$ to determine the probability of assigning treatment $a=1$; $\nu\in[0,0.5]$ sets a lower bound on the treatment randomization probability, ensuring that each arm is chosen with probability at least $\nu$; $b>0$ controls the range within which these probabilities respond smoothly to the estimated CATE $B_{n}(w_{i})$. When $B_{n}(w_{i})>b$, indicating a substantially higher expected outcome under treatment $a=1$ compared with $a=0$, the treatment allocation probability of $a=1$ reaches the maximum $1-\nu$; when $B_{n}(w_{i})<-b$, the treatment allocation probability of $a=1$ drops to its minimum value $\nu$. When $B_{n}(w_{i})\in[-b,b]$, that probability varies smoothly with the value of CATE. As the estimated CATE approaches to 0.5, the probability gets closer to 0, balancing exploration and exploitation. In simulation, we fix $b=1$ and set $\nu_{i}=exp(-t_{i})$ for every unit $i$, where $t_{i}=\left\lceil\frac{i-n_{0}}{250}\right\rceil$ with $n_{0}=1000$. This ensures that the chance of assigning estimated optimal personalized treatment grows as the experiment progresses.

### v2

@@ 8 Application and Generalization
8.1 Application
Our proposed framework is directly applicable for cases with independent data structure. For example, consider estimating causal estimands from data collected across multiple sites where the site index does not affect the primary outcome but introduces heterogeneity in treatment distribution across sites. Specifically, define the full data of interest as $X=(W,Y(a))\sim P^{F}\in\mathcal{M}^{F}$, where $W$ denotes the baseline covariates; $Y(a)$ is the counterfactual outcome with treatment level $a\in\mathcal{A}$. Let the average treatment effect $\Psi^{F}(P^{F})=E_{P^{F}}[Y(1)-Y(0)]$ be the causal estimand of interest. Suppose one observe $n$ independent observations $O_{1},\cdots,O_{n}$ with $O_{i}=(S_{i},W_{i},A_{i},Y_{i})$, where $W_{i}\sim Q_{W}$ and $Y_{i}\sim Q_{Y}(\cdot|A_{i},W_{i})$ for $Q_{W}\in\mathcal{Q}_{W}$ and $Q_{Y}\in\mathcal{Q}_{Y}$. For every unit $i$, $S_{i}\in\{1,\dots,J\}$ is the site index that determines the treatment distribution of $A_{i}\sim g_{i}(\cdot|W_{i})$ with $g_{i}=g_{s_{i}}$. Each site $j\in[J]$ has its own treatment assignment mechanism $g_{j}:\mathcal{W}\times\mathcal{A}\to[0,1]$. Note that each site does not impact the outcome given treatment and baseline covariates, and $\Psi^{F}(P^{F})$ can be identified by the target estimand $\psi(Q)=E[\bar{Q}(1,W)-\bar{Q}(0,W)]$ under ignorability and positivity assumptions, where $\bar{Q}(A,W):=E[Y|A,W]$.
Define $\bar{g}_{n}:=\frac{1}{n}\sum_{i=1}^{n}g_{i}$. With the convergence of $\bar{g}_{n}$ to a fixed $\bar{g}_{\infty}$ (e.g. $\bar{g}_{\infty}=\sum_{j\in J}p_{s}(j)g_{j}$ for a probability mass function $p_{s}$ of the site index), one can construct an ADL-TMLE by solving the score equation $\frac{1}{n}\sum_{i=1}^{n}\frac{2A_{i}-1}{\bar{g}_{n}(A_{i}|W_{i})}[Y_{i}-\bar{Q}_{n}^{*}(A_{i},W_{i})]=0$. This yields an asymptotically normal and efficient estimator that inherits the advantages of ADL-TMLE established for adaptive design settings. This approach also applies to other settings where the site index $S$ is replaced by other variables that do not directly affect the outcome given treatment and other baseline covariates, but only influence the treatment assignment probabilities.
8.2 The General Framework
We present the general estimation and design framework for various causal estimands in broader settings, including those with longitudinal data structure.
Data, Model and Target Estimand. Consider a longitudinal data structure $O=\left(L(0),A(0),L(1),A(1),\cdots,L(K),A(K),L(K+1)\equiv Y\right)$. Let
@@ 8.2 The General Framework
be a model with a known $g$, where $q(o)=\prod_{t=0}^{K+1}q_{t}(o)$ and $g(o)=\prod_{t=0}^{k}g_{t}(o)$, which corresponds to an i.i.d. likelihood with each $L(t)\sim q_{t}(\cdot|Pa(L(t)))$ and $A(t)\sim g_{t}(\cdot|Pa(A(t)))$, where $Pa(L(t))$ and $Pa(A(t))$ denote the set of nodes that appear before $L(t)$ and $A(t)$ in the longitudinal structure of $O$, respectively. Let $g^{*}\in\mathcal{G}^{*}$ denote a fixed function that defines the causal estimand of interest as $\Psi^{F}(Q)=EY_{g^{*}}=\int_{o}y\,q(o)\,g^{*}(o)\,d\mu(o)$. This parameter only depends on the $Q$-component of the likelihood.
@@
$p^{n}_{q,\mathbf{g}_{1:n}}(\bar{o}(n))=\prod_{i=1}^{n}q(o_{i})\,g_{i}(o_{i})=\prod_{i=1}^{n}\prod_{t=0}^{K+1}q_{t}(o_{i})\prod_{t=0}^{K}g_{i,t}(o_{i}).$
@@
Semiparametrically Efficient Estimation. Define $\bar{g}_{n}=\frac{1}{n}\sum_{i=1}^{n}g_{i}$ as the average of the whole product $g_{i}=\prod_{t}g_{i,t}$. Let $D(Q,g)(O_{i})$ denote the efficient influence curve in the i.i.d. model $\mathcal{M}(g)$. To estimate $EY_{g^{*}}$ from data generated by the model $\mathcal{M}^{n}$, it is useful to consider a reference i.i.d. model $\mathcal{M}(\bar{g}_{n})$ with the average design $\bar{g}_{n}$. In that i.i.d. model, the efficient influence curve for the parameter $\Psi(Q)$ is given by $D(Q,\bar{g}_{n})(O)$. Importantly, any TMLE constructed by solving $n^{-1}\sum_{i=1}^{n}D(Q,\bar{g}_{n})(O_{i})=0$ under the i.i.d. model with $\bar{g}_{n}$ is also an efficient estimator for the original model $\mathcal{M}^{n}$ for the observed data, with its asymptotic normality supported by Martingale theories and its semiparametric efficiency provided by the convolution theorem, as shown in Sections 4 and 5. This unified estimation framework applies not only to adaptive design settings discussed above, but also to other contexts with longitudinal data structure with dynamic interventions (Robins 1986, van der Laan & Gruber 2012) and time-series structures (van der Laan & Malenica 2018, Malenica et al. 2021). For the design problem, the construction of the self-calibrating $\bar{g}$-driven adaptive design implied by this efficient estimation approach directly mirrors that in Section 6.
@@ 9 Conclusions
In this work, we propose a novel ADL-TMLE framework for estimating a wide class of causal estimands from sequential adaptive experiments. We establish both asymptotic normality and semiparametric efficiency of ADL-TMLE, and demonstrate superior variance-reduction performance in finite samples. Notably, our framework substantially relaxes the requirements for adaptive experiments that enable valid statistical inference, as it only requires positivity and stabilization for the average design. Motivated by these results, we further propose a novel $\bar{g}$-driven adaptive design that dynamically steers the average design toward the estimated optimal design in a stable self-calibrating manner. This adaptive design achieves faster convergence to the oracle design and yields improved estimation efficiency compared with standard approach. Finally, we provide a general estimation and design framework applicable to broader settings with longitudinal structures. Future research could explore extensions to other related settings and broader areas of applications.
@@
Since ADL-TMLE solves $\frac{1}{n}\sum_{i=1}^{n}D(Q_{n}^{*},\bar{g}_{n})(O_{i})=0$, we have
$\psi(Q_{n}^{*})-\psi(Q_{0})=\frac{1}{n}\sum_{i=1}^{n}[D(Q_{n}^{*},\bar{g}_{n})(O_{i})-P_{Q_{0},\bar{g}_{n}}D(Q_{n}^{*},\bar{g}_{n})].$
By adding and subtracting $D(Q_{\infty},\bar{g}_{n})(O_{i})$, $D(Q_{\infty},\bar{g}_{\infty})(O_{i})$, and their conditional expectations given $\bar{O}(i-1)$, we obtain the desired decomposition equation. ∎
@@
$=$ $\Delta_{n,h}-\frac{1}{2}\|h\|_{\mathcal{H}(Q,\bar{g}_{\infty})}^{2},$
where we denote $\Delta_{n,h}:=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}h(O_{i})+o_{P^{n}_{Q}}(1)$, with the first term a zero-mean martingale. The second equation is based on the convergence of conditional variance of the zero-mean martingale according to Theorem 2.23 of Hall & Heyde (2014), as $E[h^{2}(O_{i})|\bar{O}(i-1)]$ is bounded. The third equation is based on the dependence of $g_{i}$ on the past $\bar{O}(i-1)$ in our adaptive design setup and the definition of $\|h\|_{\mathcal{H}(Q,\bar{g}_{n})}^{2}$. The last equation relies on the convergence of the adaptive design described in Assumption 10 and results in $\Delta_{n,h}\overset{d}{\to}\Delta_{h}\sim N(0,\|h\|_{\mathcal{H}(Q,\bar{g}_{\infty})}^{2})$ by Martingale central limit theorem. ∎
@@
With Assumption 10, the log likelihood ratio for the perturbed distribution of $P^{n}_{Q_{0}}$ is obtained by $log\frac{dP^{n}_{Q_{\epsilon_{n}}^{h}}(O_{1},\cdots,O_{n})}{dP^{n}_{Q_{0}}(O_{1},\cdots,O_{n})}=\Delta_{n,h}-\frac{1}{2}\|h\|_{\mathcal{H}(Q_{0},\bar{g}_{\infty})}^{2}$, where $h\in\mathcal{H}(Q_{0},\bar{g}_{\infty})$ and $\Delta_{n,h}\overset{d}{\to}\Delta_{h}\sim N(0,\|h\|_{\mathcal{H}(Q_{0},\bar{g}_{\infty})}^{2})$ under $P^{n}_{Q_{0}}$ via Theorem 7. Then, by Le Cam’s third lemma (Theorem 3.10.7 in van der Vaart et al. (1996)), we have that
@@
By the convolution theorem (Theorem 8), the asymptotic variance of $\sqrt{n}(T_{n}-\psi(Q_{0}))$ for regular estimators $T_{n}$ is no less than $\|\dot{\psi}^{*}(Q_{0})\|_{\mathcal{H}(Q_{0},\bar{g}_{\infty})}^{2}$, which equals $P_{Q_{0},\bar{g}_{\infty}}D(Q_{0},\bar{g}_{\infty})^{2}$ and is achieved by the asymptotic variance of $\psi(Q_{n}^{*})$. Therefore, $\psi(Q_{n}^{*})$ is efficient. ∎
@@
The data generating distribution is as follows: $\bar{Q}_{0}(1,W)=25+10/\left[1+exp(-(2W+1))\right]$; $\bar{Q}_{0}(0,W)=20+17.5/\left[1+exp(-(W+0.1))\right]$; $Var_{Q_{0}}[Y|A=1,W]=1+3W^{3}$; $Var_{Q_{0}}[Y|A=0,W]=1+1.5\times(3-W)^{3}.$ Figure 5 shows the true conditional expectation and variance of the outcome given different treatment levels and baseline covariates.
For benefit-driven design, we guide treatment randomization probabilities toward optimal personalized treatment using baseline covariates through estimating the conditional average treatment effect (CATE). Define the true CATE function as $B_{0}:\mathcal{W}\to\mathbb{R}$, with $B_{0}(W)=\bar{Q}_{0}(1,W)-\bar{Q}_{0}(0,W)$. For a newly enrolled unit $i$ with baseline covariates $w_{i}$, we use $B_{n}(w_{i})$ to denote the estimate of $B_{0}(w_{i})$ based on the current data $\bar{O}(i-1)$. We estimate CATE functions by first transforming the outcomes into doubly robust pseudo-outcomes (Rubin & van der Laan 2005, van der Laan & Rubin 2006, van der Laan & Luedtke 2015, Kennedy 2023), then regressing them on baseline covariates using Super Learner (van der Laan et al. 2007). This estimated CATE guides the treatment randomization probabilities toward the estimated optimal personalized treatment $I(B_{n}(w_{i})>0)$ through a function $\Gamma_{\nu,b}$ introduced by Chambaz et al. (2017):
@@
Here, $B_{n}(w_{i})$ is used as input to $\Gamma_{\nu,b}$ to determine the probability of assigning treatment $a=1$; $\nu\in[0,0.5]$ sets a lower bound on the treatment randomization probability, ensuring that each arm is chosen with probability at least $\nu$; $b>0$ controls the range within which these probabilities respond smoothly to the estimated CATE $B_{n}(w_{i})$. When $B_{n}(w_{i})>b$, indicating a substantially higher expected outcome under treatment $a=1$ compared with $a=0$, the treatment allocation probability of $a=1$ reaches the maximum $1-\nu$; when $B_{n}(w_{i})<-b$, the treatment allocation probability of $a=1$ drops to its minimum value $\nu$. When $B_{n}(w_{i})\in[-b,b]$, that probability varies smoothly with the value of CATE. As the estimated CATE approaches to 0.5, the probability gets closer to 0, balancing exploration and exploitation. In simulation, we set $b=1$ and $\nu_{i}=exp(-t_{i})$ for every unit $i$, where $t_{i}=\left\lceil\frac{i-n_{0}}{250}\right\rceil$ with $n_{0}=1000$, thereby increasing the chance of assigning substantially better treatments as the experiment progresses.